OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 9 and 36. What is the length of the altitude? 6 18 45 3

OpenStudy (anonymous):

@Ishaan94

OpenStudy (anonymous):

l^2 + 9^2 = b^2 l^2 + 36^2 = p^2 2l^2 + 36^2 + 9^2 = 45^2 solve the equation for l.